OpenStudy (anonymous):

Which of the numbers for n makes the following equation TRUE? 3(2n − 1) − 2n = 21 A. 4 B. 5 C. 6 D. 7

OpenStudy (anonymous):

Hi...can you help me??????????

OpenStudy (solomonzelman):

I want to be more creative. \(\Large\color{black}{ \bf 3(2n − 1) − 2n = 21 }\) \(\Large\color{black}{ \bf 3(2n − 1) − 2n +1= 21 +1 }\) \(\Large\color{black}{ \bf 3(2n − 1) − (2n -1)= 21 +1 }\) \(\Large\color{black}{ \bf 3(2n − 1) − (2n -1)= 22 }\) \(\Large\color{black}{ \bf 3(2n − 1) − 1(2n -1)= 22 }\) \(\Large\color{black}{ \bf 2(2n -1)= 22 }\) \(\Large\color{black}{ \bf (2n -1)=11 }\) \(\Large\color{black}{ \bf 2n -1=11 }\) \(\Large\color{black}{ \bf 2n -1+1=11 +1}\) \(\Large\color{black}{ \bf 2n =12}\) and you know the final step..

OpenStudy (anonymous):

Divison

OpenStudy (solomonzelman):

yes

OpenStudy (anonymous):

Would it be 6

OpenStudy (solomonzelman):

yes

OpenStudy (anonymous):

Which of the numbers for n makes the following equation TRUE? 4(n + 3) + 7n = 34 A. −2 B. 0 C. 2 D. 4

OpenStudy (solomonzelman):

For this equation I am just going to provide you with the steps. 1) expand 4(n+3) 2) add like terms on the left side 3) subtract 12 from both sides 4) divide both sides by 11 there is your answer.

OpenStudy (anonymous):

Ok.....Thx

OpenStudy (anonymous):

I got 4?

OpenStudy (solomonzelman):

No, not 4.

OpenStudy (anonymous):

I think I did my math wrong let me check again

OpenStudy (solomonzelman):

you can post the work in here, if you want :)

OpenStudy (anonymous):

I got two choices its either 0 or -2

OpenStudy (solomonzelman):

again not right.

OpenStudy (anonymous):

So the only thing to choose from is positive 2

OpenStudy (solomonzelman):

yes, that's the only correct answer.

OpenStudy (anonymous):

Do you mind helping me with more or do you have to help someone else???

OpenStudy (solomonzelman):

I have to do something else right now, online. Not openstudy-related.

OpenStudy (anonymous):

Can you help me Wolf1728?????

OpenStudy (anonymous):

Solve the following linear equation for x. 2x + 7 = 2(6 + x) – 4 A. x = -2 B. x = 5/2 C. infinitely many solutions D. no solution

OpenStudy (anonymous):

Thats my question...That I just posted

OpenStudy (anonymous):

Solve the following linear equation for x. 2x + 7 = 2(6 + x) – 4 A. x = -2 B. x =5/2 C. infinitely many solutions D. no solution

OpenStudy (solomonzelman):

\(\Large\color{blue}{ \bf 2x + 7 = 2(6 + x) – 4 }\) \(\Large\color{blue}{ \bf 2x + 7 =12 + 2x– 4 }\) \(\Large\color{blue}{ \bf 2x + 7 = 2x+8 }\) subtract 2x from both sides and you get, \(\Large\color{blue}{ \bf 7 = 8 }\)

OpenStudy (solomonzelman):

So what do you think ?

OpenStudy (anonymous):

No solution??

OpenStudy (solomonzelman):

Yes, correct. No solution.

OpenStudy (anonymous):

What are the solutions, if any, of the following equation? 6y − 3 = −6y + 3 A. There is no solution. B. There are an infinite number of solutions. C. 0.5 D. 2

OpenStudy (anonymous):

ind the solution of the linear equation. 4a - 11 = 2a + 2(a + 6) A. a = 1 B. a = 12/11 C. infinitely many solutions D. no solution Thats the the correct question Sorry

OpenStudy (anonymous):

i think it is D i am not to sure tho sorry

OpenStudy (anonymous):

Thats ok at least you tried your best